OpenStudy (anonymous):

Please help me solve? I'm looking at this problem and not seeing how to work it so that my answer comes out looking like the choices...

OpenStudy (anonymous):

\[2\sqrt{44x ^{3}}-\sqrt{7}-\sqrt{99x ^{3}}+\sqrt{63}\]

OpenStudy (anonymous):

Combine like terms, right? I don't see how to combine \[2\sqrt{44x ^{3}}\] and \[-\sqrt{99x ^{3}}\]

OpenStudy (anonymous):

before combining see if you can simplify each term first

OpenStudy (anonymous):

How so...?

OpenStudy (anonymous):

Oh! like 11 goes into both 44 and 99??

OpenStudy (anonymous):

see if u can get rid of the roots!

OpenStudy (anonymous):

How would I go about doing that? Right now I have \[2\sqrt{4x ^{3}}-\sqrt{7}-\sqrt{9x ^{3}}+\sqrt{63}\]

OpenStudy (anonymous):

Does it look right? Going in the right direction maybe?

OpenStudy (anonymous):

Oh, and I eliminated 7. and the sqrt of 63 is now the sqrt of 9

hartnn (hartnn):

\(\sqrt {63} = \sqrt {9} \sqrt {7}=...?\)

OpenStudy (anonymous):

Now it looks like \[2\sqrt{4x ^{3}}-\sqrt{9x ^{3}}+\sqrt{9}\]

hartnn (hartnn):

\(\sqrt {44x^3} = \sqrt {4} \sqrt{x^2} \sqrt{11x}=...?\) you just can't eliminate 7...

OpenStudy (anonymous):

something is missing

hartnn (hartnn):

\(\sqrt {44x^3} = \sqrt {4} \sqrt{x^2} \sqrt{11x}=...? \\\sqrt {99x^3} = \sqrt {9} \sqrt{x^2} \sqrt{11x}=...?\)

OpenStudy (anonymous):

But & goes into 7 one time and it goes into 63 9 times. Simplifying... Right?

OpenStudy (anonymous):

but 7*

hartnn (hartnn):

so, \(\sqrt {63}-\sqrt 7 = \sqrt{9}\sqrt{7}-\sqrt{7} = \sqrt{7}[3-1]=..?\)

hartnn (hartnn):

you see what i did there ^ ?

OpenStudy (anonymous):

No... :(

OpenStudy (anonymous):

@hartnn i see what he did

OpenStudy (anonymous):

3x3

hartnn (hartnn):

\(\sqrt {xy}=\sqrt{x}\sqrt{y}\) ok ? so, \(\sqrt{63}=\sqrt{9}\sqrt{7}=3 \sqrt{7}\) right ?

hartnn (hartnn):

ask if any doubts.

OpenStudy (anonymous):

That does not look like any of my answer choices :(

hartnn (hartnn):

that was just simplification of constant terms.... \(3\sqrt 7- \sqrt 7= 2 \sqrt 7\) simplification of 'x' terms is still remaining. and more important thing is that you understand....so that you can do other problems on your own...

OpenStudy (anonymous):

I have no idea.

hartnn (hartnn):

so, do you want me to start over ? or take it from a particular step ?

OpenStudy (anonymous):

Start from the beginning please?

hartnn (hartnn):

right, lets see term by term, so 63 is 9 times 7 so, \(\sqrt {63}= \sqrt {9}\sqrt{7}=...?\)

OpenStudy (anonymous):

Yes. And what do we do with the \[\sqrt{7}\] just leave it where it is?

hartnn (hartnn):

i was getting there... you first tell me what will be \(\sqrt 9 \sqrt7 =... ?\)

OpenStudy (anonymous):

simplify by dividing both \[\sqrt{7}\] and \[\sqrt{63}\] by 7?

hartnn (hartnn):

dividing ? why ? no..

OpenStudy (anonymous):

Oh... Grrr

hartnn (hartnn):

lets finish 1 term entirely \(\sqrt {63}=\sqrt 9 \sqrt7 =... ?\)

OpenStudy (anonymous):

Yeah. 3.

hartnn (hartnn):

so, \(\sqrt {63}=\sqrt 9 \sqrt7 =... ?\)

OpenStudy (anonymous):

\[\sqrt{3}\] \[\sqrt{2.6}??\]

hartnn (hartnn):

since \(\sqrt 9=3\) \(\sqrt {63}=\sqrt 9 \sqrt7 =3 \sqrt 7\) got this ?

OpenStudy (mathstudent55):